# An interpretation of the Brownian bridge as a physics-informed prior for the Poisson equation

### v1

@@ 3 The Brownian bridge as a physics-informed prior
We establish an explicit connection between the Brownian bridge GP estimator and the physics-informed loss function. In particular, we show the posterior mean function when starting with a shifted Brownian bridge GP is exactly the function which minimizes the physics-informed loss of eq. (1.4), under certain criteria. We demonstrate this for the case where the posterior remains Gaussian, and also when the posterior is non-Gaussian, as a MAP estimator in 1D. We begin with the simpler case where we have point measurements with additive Gaussian noise according to eq. (2.3).
@@ 3.1 Physics-informed prior as a Gaussian process
Let $L$ denote the minus Laplacian operator on $L^{2}(\Omega)$, i.e., $(Lu)(x)=-\nabla^{2}u|_{x}$. We will see later that $L$ is the precision operator associated with the GP we are after. Denote the inverse of $L$ by $C$. This is the operator with kernel given by the Green’s function of the Laplacian, i.e., $C$ is the operator defined by
[…]
In our example where $\Omega=[0,1]^{d}$, the covariance kernel $k$ is best expressed with the Mercer representation. One can check the orthonormal eigenfunctions associated with $C$ are
$\psi_{n_{1},\dots,n_{d}}(x)=2^{d/2}\sin(n_{1}\pi x)\cdots\sin(n_{d}\pi x),$ (3.2)
@@
$\lambda_{n_{1}\dots n_{d}}=\frac{1}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}.$ (3.3)
@@
$k(x_{1},\dots,x_{d},x_{1}^{\prime},\dots,x_{d}^{\prime})=2^{d}\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\frac{\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}.$ (3.4)
Note this kernel is associated with the Brownian bridge in $d$-dimensions. For example, on $[0,1]$, the kernel is simply $k(x,x^{\prime})=\min\{x,x^{\prime}\}-xx^{\prime}$, which is exactly the covariance kernel of the Brownian bridge process. Typical sample path behavior of the unit Brownian bridge in 1D can be seen in Figure 1.
Figure 1: Typical unit Brownian bridge behavior. Sample paths are shown in black, and the shaded region represents two standard deviations. The variance is given by $\mathbb{V}[u(x)]=x(1-x)$.
We now show that the physics-informed loss function emits a KRR objective with this kernel.
@@
The physics-informed loss function given by eq. (1.4) is equivalent to shifted kernel ridge regression objective
$\mathcal{L}(u)=\frac{1}{n}\sum_{i=1}^{n}(u(x_{i})-y_{i})^{2}+\frac{\eta}{2}\|u-Cq\|_{H_{k}}^{2},$ (3.5)
@@
[…]
@@
[…]
$=\frac{1}{2}\langle u-Cq,L(u-Cq)\rangle,$
@@
$(Lu)(x)$ $=\int_{\Omega}\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1}\dots,n_{d}}^{-1}\psi_{n_{1},\dots,n_{d}}(x)\psi_{n_{1},\dots,n_{d}}(x^{\prime})u(x^{\prime})dx^{\prime}$
$=\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1}\dots,n_{d}}^{-1}\psi_{n_{1},\dots,n_{d}}(x)\langle u,\psi_{n_{1},\dots,n_{d}}\rangle.$
Plugging this into the quadratic form, we get
$E(u)$ $=\frac{1}{2}\left\langle u-Cq,\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1}\dots,n_{d}}^{-1}\psi_{n_{1},\dots,n_{d}}\langle u-Cq,\psi_{n_{1},\dots,n_{d}}\rangle\right\rangle$
$=\frac{1}{2}\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1}\dots,n_{d}}^{-1}\left\langle u-Cq,\psi_{n_{1},\dots,n_{d}}\langle u-Cq,\psi_{n_{1},\dots,n_{d}}\rangle\right\rangle$
$=\frac{1}{2}\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1}\dots,n_{d}}^{-1}\langle u-Cq,\psi_{n_{1},\dots,n_{d}}\rangle^{2},$
$=\frac{1}{2}\|u-Cq\|_{H_{k}}^{2}$
where the last line holds by Mercer representation of the RKHS-norm. ∎
So, we have connected the energy-based physics-informed loss function to KRR. The loss function which is used to train physics-informed models, such as PINNs, for the Poisson equation is exactly the objective function in KRR with the Brownian bridge kernel. Observe that the objective function is shifted to minimize the distance in RKHS between the estimator and $Cq$. As $C$ is defined by the Green’s function, $Cq$ is the unique solution to eq. (1.1), and we verify that the physics-informed model is indeed trying to find the closest possible match to the solution of eq. (1.1), while also fitting the data. In fact, unlike the integrated square residual which seeks to minimize the $L^{2}(\Omega)$-norm, this objective function is also trying to match the smoothness. This is evidenced later when we identify $H_{k}\subset H^{1}(\Omega)$.
From Proposition 3.1, it is now fairly trivial to connect the physics-informed loss function to a GP regression scheme. Recall now Proposition 2.1, which shows an equivalence between the GP posterior mean function and the KRR estimator. That is, $\tilde{m}$ is the function which solves the problem
@@
Inspired by this, we define the GP prior, which we term as the physics-informed prior for the Poisson equation, $u\sim\mathcal{GP}(u^{0},\beta^{-1}k)$, $u^{0}=Cq$, and $k$ is the Brownian bridge kernel as given by eq. (3.4). We have included a hyperparameter $\beta\in(0,\infty)$ in order to control the variance of this prior. Later we will prove that $\beta$ plays a key role in detecting model-form error. Notice that the prior is centered at the unique solution to eq. (1.1).
The prior allows the sample paths to vary around $u^{0}$, which is desirable in the case of an imperfect model. For the Brownian bridge, the variance reaches its maximum in the center of the domain, with no variance on $\partial\Omega$. The additional hyperparameter $\beta$ controls the magnitude of the variance, which can be seen in the limiting cases. As $\beta\to 0$, $\mathbb{V}[u]\to\infty$. This essentially corresponds to placing a flat, uninformative prior on $L^{2}(\Omega)$, and the physics play no role. If $\beta\to\infty$, the prior collapses to a Dirac centered at $u^{0}$. This corresponds to the ultimate belief that the underlying field truly is governed by the Poisson equation. The only field we consider is the one a priori assumed to be correct. It is for this reason that we view the prior as a soft-constraint for the physics, with $\beta$ encoding the degree of model-trust. An example of this behavior is shown in Figure 2.
@@
Setting $\eta=\sigma^{2}\beta/n$ gives an equivalence between the posterior mean function, when starting with the physics-informed prior, and the minimizer of the physics-informed loss function. We summarize this result in the following theorem.
@@
Consider training data of the form $y_{i}=u(x_{i})+\gamma_{i}$, $i=1,\dots,n$, where $\gamma_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2})$ and let $u:\Omega\to\mathbb{R}$ be the target function. Let $E$ be the energy functional for the Poisson equation, i.e. $E(u)=\int\frac{1}{2}\|\nabla u\|^{2}+qu\>d\Omega$. Letting $\eta=\sigma^{2}\beta/n$, we have $\tilde{m}=\hat{u}$, where
  1. (i)
@@
  2. (ii)
    $\hat{u}$ is the solution to the physics-informed optimization problem
    $\hat{u}=\operatorname*{arg\,min}_{u\in H_{k}}\>\mathcal{L}_{\mathrm{data}}(u)+\eta E(u).$
The above theorem allows us to analyze the behavior of the physics-informed machine learning approach through the established theory of GP regression. For example, we can study convergence conditions for the field reconstruction problem and for inverse problems. This is reserved for later sections.
@@ 3.2 Physics-informed prior as a Gaussian measure
We derive a similar result to Theorem 3.1 in the setting of infinite-dimensional Bayesian inverse problems [44] in the 1D case. This is useful, for instance, in the case where the measurement operator is nonlinear, and we cannot easily rely on the GP formulae. We use this relationship to derive a physics-informed loss function, modified to account for a general measurement operator, by identifying the functional with produces the MAP estimate of the inverse problem starting with the Brownian bridge as the prior measure.
The reason we must restrict ourselves to 1D is the fact that the covariance operator associated with the Brownian bridge kernel is trace class only when $d=1$. That is, $\mathrm{tr}\left(C\right)=\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1},\dots,n_{d}}<\infty$ for $d=1$ and diverges for $d>1$. The GP prior is related to a Gaussian measure on $L^{2}(\Omega)$ in the following manner
Theorem 3.2 (Theorem 2 [41]).
@@
[…]
For derivations, it is often convenient to shift the space so that the prior is centered. According to Theorem A.3, this is permitted so long as $u^{0}\in H_{k}$. As $u^{0}$ is exactly the solution of eq. (1.1), we have $u^{0}\in H_{p}$. Later in Lemma 4.1, we show that
$H_{k}=\{u\in H^{1}(\Omega):u=0\>\mathrm{on}\>\partial\Omega\}.$
Then, $u^{0}\in H_{p}\subset H_{k}$, and the shift is justified.
Let $\mathcal{X}$ denote the function space for which the target function lives. Suppose now we have data $d\in\mathbb{R}^{n}$ generated according to $y=R(u)+\gamma$, where $R:\mathcal{X}\to\mathbb{R}^{n}$ is the observation map, which is in general nonlinear, and $\gamma\sim\mathcal{N}(0,\Gamma)$ is an additive noise process. Following the Bayesian approach [44], we look to derive the posterior in function space. To identify the posterior measure $\mu^{y}$, we apply Bayes’s rule, which takes the following form in infinite-dimensions.
Theorem 3.3 (Bayes’s theorem [28, 44]).
@@
Theorem 3.3 admits a closed form expression in a special case. Assuming that $R$ is linear, the posterior $\mu^{d}$ is also Gaussian $\mathcal{N}(\tilde{m},\tilde{C})$, with
@@
There is a minor technicality to discuss here about the existence and interpretation of $\mu^{y}$ as a posterior measure. The prior measure must be chosen such that $\mu_{0}(\mathcal{X})=1$. In much of the literature, the measurement operator involves solving a PDE, in which case care must be taken when choosing the prior. The advantage of our approach is that the physics are encoded into the prior, rather than the likelihood. For the Brownian bridge, $\mu_{0}(L^{2}(\Omega))=1$, so the only requirement is that $R$ acts on $L^{2}(\Omega)$, a fairly trivial assumption.
To identify the MAP estimate of the posterior, we follow the work laid out in [17]. The precise definition is as follows.
Definition 3.1 (MAP estimate of a Gaussian measure).
Let $\mu\sim\mathcal{N}(0,C)$ be a Gaussian measure on a separable Banach space $\mathcal{X}$, and assume the posterior distribution $\mu^{y}$ has density with respect to $\mu_{0}$ given by
$\frac{d\mu^{y}}{d\mu_{0}}(u)\propto\exp\left\{-\Phi(u)\right\}$
For $u\in\mathcal{X}$, denote the open ball centered at $u$ with radius $\delta>0$ by $B(u;\delta)\subset\mathcal{X}$. For fixed $\delta$, let $u^{\delta}=\operatorname*{arg\,max}_{u\in\mathcal{X}}\mu^{y}(B(u;\delta))$. A point $\tilde{u}\in\mathcal{X}$ satisfying
$\lim_{\delta\to 0}\frac{\mu^{y}(B(\tilde{u};\delta))}{\mu^{y}(B(u^{\delta};\delta))}=1$
is a MAP estimate for $\mu$.
The MAP estimate may be identified through the Onsager-Machlup functional [19, 22]. This is the functional $I:H_{k}\to\mathbb{R}$ such that
$\lim_{\delta\to 0}\frac{\mu^{y}(B(u_{2};\delta))}{\mu^{y}(B(u_{1};\delta))}=\exp\{I(u_{1})-I(u_{2})\},$
where $B(u_{i};\delta)$ is the open ball in $L^{2}(\Omega)$ centered at $u_{i}$ with radius $\delta$. For fixed $u_{1}$, any function $u_{2}$ which minimizes the Onsager-Machlup functional can be taken as the MAP estimate. For our specific problem, the Onsager-Machlup functional is
$I(u)=\left\{\begin{split}&\Phi(u)+\frac{1}{2}\|u-u^{0}\|^{2}_{H_{k}},\quad\mathrm{if}\>u-u^{0}\in H_{k}\\ &+\infty,\quad\mathrm{otherwise},\end{split}\right.$ (3.6)
as show in [17, Theorem 3.2]. So, any MAP estimate of $\mu^{y}$ will live in the Cameron-Martin space (which is also $H_{k}$) of $\mu_{0}$. Further, if $\Phi$ is linear in $u$ this MAP estimate is unique.
Equation 3.6 is a natural candidate to build the physics-informed loss function for a general measurement operator. Adjusting the notation a bit to match a loss function, the MAP estimate of the Gaussian measure solves the problem
$\hat{u}=\operatorname*{arg\,min}_{u\in H_{k}}\frac{1}{2}\|\Gamma^{-1}(y-R(u))\|^{2}+\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}+qu\>d\Omega,$ (3.7)
which follows from Proposition 3.1. If the data are collected according to eq. (2.3), then it is easy to verify that eq. (3.7) reduces to the original physics-informed loss function.

### v2

@@ 3 The Brownian bridge as a physics-informed prior
We establish an explicit connection between the Brownian bridge GP estimator and the physics-regularized inverse problem. In particular, we show the posterior mean function when starting with a shifted Brownian bridge GP is exactly the function which minimizes the variational problem eq. (1.4), under certain criteria. We demonstrate this for the case where the posterior remains Gaussian, and also when the posterior is non-Gaussian, as a MAP estimator in 1D. We begin with the simpler case where we have point measurements with additive Gaussian noise according to eq. (2.3).
Before proceeding, we discuss the regularity conditions so that the definitions and operations that follow are well-defined. Because we are modeling the solution to eq. (1.1) as a GP, we would ideally like to be able to evaluate it pointwise. So, at the very least, we ask that the solution also belongs to the space of bounded, continuous functions over the domain, which we will denote by $\mathcal{C}^{0}_{B}(\Omega).$ The required conditions are given by the next result.
Proposition 3.1.
Assume that eq. (1.1) possesses a weak solution $u^{0}\in H_{0}^{1}(\Omega)$. Further assume that if $d\leq 3$, then $q\in L^{2}(\Omega)$, and for $d>3$, $q\in H^{\tau}(\Omega)$ with $\tau>d/2-2$. Then $u^{0}\in H^{2}(\Omega)$ for $d\leq 3$ and $u^{0}\in H^{\tau+2}(\Omega)$ for $d>3$. In both cases, there exists a version of $u^{0}$ that is equivalent to $u^{0}$ a.e., denoted by $u^{\dagger}$, with $u^{\dagger}\in C^{0}_{B}(\Omega)$.
Proof.
We show that $u^{0}$ lives in the stated Sobolev spaces, which are then embedded in $C^{0}_{B}(\Omega)$. The case $d\leq 3$ follows from [evans2022partial, Sec. 6.3, Theorem 4], and $d>3$ results from [evans2022partial, Sec. 6.3, Theorem 5]. The continuity follows from a case of the Sobolev embedding theorem [adams2003sobolev, Theorem 4.12], which states that for integers $s\geq 0$, $\tau\geq 1$, $H^{s+\tau}(\Omega)$ is embedded in $C^{s}_{B}(\Omega)$ when $2\tau>d$. The result follows in each case by setting $s=0$. ∎
Moving forward, we will assume that Proposition 3.1 holds, and by the equivalence relation given by the embedding, we will take the weak solution to be $u^{\dagger}$ so that we may evaluate it pointwise. So that there is no confusion, we will still refer to this function as $u^{0}$.
@@ 3.1 Physics-informed prior as a Gaussian process
Let $L$ denote the minus Laplacian operator on $H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$, i.e., $(Lu)(x)=-\nabla^{2}u|_{x}$. We will see later that $L$ is the precision operator associated with the GP we are after. Denote the inverse of $L$ by $C$. This is the operator with kernel given by the Green’s function of the Laplacian, i.e., $C$ is the operator defined by
$(Cu)(x)\coloneqq(L^{-1}u)(x)=\int_{\Omega}k(x,x^{\prime})u(x^{\prime})dx^{\prime},\quad u\in L^{2}(\Omega).$
In our example where $\Omega=[0,1]^{d}$, the covariance kernel $k$ is best expressed with the Mercer representation.
Notation 3.1.
Before writing the covariance kernel we introduce a multi-index notation. Let $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}^{d}$ and $|\alpha|=\sum_{i=1}^{d}\alpha_{i}$. We also write for an $x=(x_{1},\dots,x_{d})\in\Omega$, $\sin(\alpha\pi x)=\prod_{i=1}^{d}\sin(\alpha_{i}\pi x_{i})$.
One can check the orthonormal eigenfunctions associated with $C$ are [polyanin2001handbook, Sec. 8.2.2-16]
$\psi_{\alpha}(x)=2^{d/2}\sin(\alpha\pi x),$ (3.2)
@@
$\lambda_{\alpha}=\frac{1}{\pi^{2}|\alpha|^{2}}.$ (3.3)
@@
$k(x,x^{\prime})=2^{d}\sum_{|\alpha|\in\mathbb{N}}\frac{\sin(\alpha\pi x)\sin(\alpha\pi x^{\prime})}{\pi^{2}|\alpha|^{2}}.$ (3.4)
This kernel is associated with the Brownian bridge in $d$-dimensions. For example, on $[0,1]$, the kernel is simply the Green’s function of eq. (1.1) $k(x,x^{\prime})=\min\{x,x^{\prime}\}-xx^{\prime}$, which is exactly the covariance kernel of the Brownian bridge process.
We now show that the physics-regularized inverse problem emits a KRR objective with this kernel.
@@
The physics-regularized inverse problem given by eq. (1.4) is equivalent to the shifted kernel ridge regression objective
$\mathcal{L}(u)=\frac{1}{n}\sum_{i=1}^{n}(u(x_{i})-y_{i})^{2}+\frac{r}{2}\|u-Cq\|_{H_{k}}^{2},$ (3.5)
@@
$\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}-qu\>d\Omega$ $=\int_{\partial\Omega}\frac{1}{2}u\nabla u\cdot\mathbf{n}\>d\Omega-\int_{\Omega}\frac{1}{2}u\nabla^{2}u\>d\Omega\color[rgb]{1,0,0}-\color[rgb]{0,0,0}\int_{\Omega}qu\>d\Omega$
@@
$=\int_{\Omega}\frac{1}{2}(u-L^{-1}q)L(u-L^{-1}q)\>d\Omega+\mathrm{const.}$
$=\frac{1}{2}\langle u-Cq,L(u-Cq)\rangle+\mathrm{const.},$
@@
$(Lu)(x)$ $=\int_{\Omega}\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\psi_{\alpha}(x)\psi_{\alpha}(x^{\prime})u(x^{\prime})dx^{\prime}$
$=\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\psi_{\alpha}(x)\langle u,\psi_{\alpha}\rangle,$
where interchanges between summation and integration are permitted by the absolute continuity of Mercer’s theorem. Plugging this into the quadratic form, we get
$E(u)$ $=\frac{1}{2}\left\langle u-Cq,\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\psi_{\alpha}\langle u-Cq,\psi_{\alpha}\rangle\right\rangle+\mathrm{const.}$
$=\frac{1}{2}\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\left\langle u-Cq,\psi_{\alpha}\langle u-Cq,\psi_{\alpha}\rangle\right\rangle+\mathrm{const.}$
$=\frac{1}{2}\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\langle u-Cq,\psi_{\alpha}\rangle^{2}+\mathrm{const.}$
$=\frac{1}{2}\|u-Cq\|_{H_{k}}^{2}+\mathrm{const.}$
where the last line holds by Mercer representation of the RKHS-norm. The additive constant may be dropped from the optimization. ∎
So, we have established a connection between KRR with the Brownian bridge kernel and the loss function coming from the variational form of the Poisson equation. Observe that the objective function is shifted to minimize the distance in RKHS between the estimator and $Cq$. As $C$ is defined by the Green’s function, $Cq$ is the unique solution to eq. (1.1), and we verify that the model is indeed trying to find the closest possible match to the solution of eq. (1.1), while also fitting the data. In fact, unlike the integrated square residual which seeks to minimize the $L^{2}(\Omega)$-norm, this objective function is also trying to match the smoothness. This is evidenced later when we identify $H_{k}=H^{1}_{0}(\Omega)$.
From Proposition 3.2, it is now fairly trivial to connect the physics-regularized inverse problem to a GP regression scheme. Recall now Proposition 2.1, which shows an equivalence between the GP posterior mean function and the KRR estimator. That is, $\tilde{m}$ is the function which solves the problem
@@
Inspired by this, we define the GP prior, which we term as the physics-informed prior for the Poisson equation, $u\sim\mathcal{GP}(u^{0},\beta^{-1}k)$, $u^{0}=Cq$, and $k$ is the Brownian bridge kernel as given by eq. (3.4). In doing so we obtain a natural Bayesian interpretation of the familiar variational formulation of the Poisson equation. We have included a hyperparameter $\beta\in(0,\infty)$ in order to control the variance of this prior. Later we will prove that $\beta$ plays a key role in detecting model-form error. Notice that the prior is centered at the unique solution to eq. (1.1).
@@
The prior allows the sample paths to vary around $u^{0}$, which is desirable in the case of an imperfect model. For the Brownian bridge, the variance reaches its maximum in the center of the domain, with no variance on $\partial\Omega$. The additional hyperparameter $\beta$ controls the magnitude of the variance, which can be seen in the limiting cases. As $\beta\to 0$, $\mathbb{V}[u]\to\infty$. This essentially corresponds to placing a flat, uninformative prior on $L^{2}(\Omega)$, and the physics plays no role. If $\beta\to\infty$, the prior collapses to a Dirac centered at $u^{0}$. This corresponds to the ultimate belief that the underlying field truly is governed by the Poisson equation. The only field we consider is the one a priori assumed to be correct. It is for this reason that we view the prior as a soft-constraint for the physics, with $\beta$ encoding the degree of model-trust. An example of this behavior is shown in Fig. 1.
Setting $r=\sigma^{2}\beta/n$ gives an equivalence between the posterior mean function, when starting with the physics-informed prior, and the minimizer of the physics-regularized inverse problem. We summarize this result in the following theorem.
@@
Consider training data of the form $y_{i}=u(x_{i})+\gamma_{i}$, $i=1,\dots,n$, where $\gamma_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2})$ and let $u:\Omega\to\mathbb{R}$ be the target function. Let $E$ be the energy functional for the Poisson equation, i.e. $E(u)=\int\frac{1}{2}\|\nabla u\|^{2}+qu\>d\Omega$. Letting $r=\sigma^{2}\beta/n$, we have $\tilde{m}=\hat{u}$, where
  1. 1.
@@
  2. 2.
    $\hat{u}$ is the solution to the physics-regularized inverse problem ^u = arg min_u H^1_0(Ω)  L_data(u) + r E(u).
The above result allows us to analyze the behavior of the physics-regularized inverse problem through the established theory of GP regression. For example, we can study convergence conditions for the field reconstruction problem and the effect of the hyperparameters. This is reserved for later sections.
@@ 3.2 Physics-informed prior as a Gaussian measure
We derive a similar result to Theorem 3.1 in the setting of infinite-dimensional Bayesian inverse problems [stuart2010inverse] in the 1D case. This is useful, for instance, in the case where the measurement operator is nonlinear, and we cannot easily rely on the GP formulae. We use this relationship to derive a parallel variational problem, modified to account for a general measurement operator, by identifying the functional with produces the MAP estimate of the inverse problem starting with the Brownian bridge as the prior measure.
The reason we must restrict ourselves to 1D is the fact that the covariance operator associated with the Brownian bridge kernel is trace class in $L^{2}(\Omega)$ only when $d=1$. That is, $\mathrm{tr}\left(C\right)=\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\lambda_{n_{1},\dots,n_{d}}<\infty$ for $d=1$ and diverges for $d>1$. The GP prior is related to a Gaussian measure on $L^{2}(\Omega)$ in the following manner
Theorem 3.2 (Theorem 2 [rajput1972gaussian]).
@@
$\int_{\Omega}m^{2}(x)dx<\infty,\quad\int_{\Omega}k(x,x)dx<\infty.$
If the above holds then $u$ induces the Gaussian measure $\mathcal{N}(m,C)$ on $L^{2}(\Omega)$, where the covariance operator is given by $\left(Cv\right)(x)\coloneqq\int_{\Omega}k(x,x^{\prime})v(x^{\prime})dx^{\prime}$, for $v\in L^{2}(\Omega)$.
In the above theorem the condition $\int_{\Omega}k(x,x)dx<\infty$ is exactly the condition that $\mathrm{tr}\left(C\right)<\infty$. So, in one-dimension we may interpret the physics-informed prior as the Gaussian measure $\mu_{0}\sim\mathcal{N}(u^{0},\beta^{-1}C)$ on $L^{2}(\Omega)$, and follow the infinite-dimensional Bayesian framework.
For derivations, it is often convenient to shift the space so that the prior is centered. According to Theorem A.3, this is permitted so long as $u^{0}\in H_{k}$. As $u^{0}$ is exactly the solution of eq. (1.2), we have $u^{0}\in H^{1}_{0}(\Omega)$. Later in Lemma 4.1, we show that $H_{k}=H^{1}_{0}(\Omega)$ so that the shift is justified.
Let $\mathcal{X}$ denote the function space for which the target function lives. Suppose now we have data $y\in\mathbb{R}^{n}$ generated according to $y=R(u)+\gamma$, where $R:\mathcal{X}\to\mathbb{R}^{n}$ is the observation map, which is in general nonlinear, and $\gamma\sim\mathcal{N}(0,\Gamma)$ is an additive noise process. Following the Bayesian approach [stuart2010inverse], we look to derive the posterior in function space. To identify the posterior measure $\mu^{y}$, we apply Bayes’s rule, which takes the following form in infinite-dimensions.
Theorem 3.3 (Bayes’s theorem [kaipio2006statistical, stuart2010inverse]).
@@
[…]
@@
There is a minor technicality to discuss here about the existence and interpretation of $\mu^{y}$ as a posterior measure. The prior measure must be chosen such that $\mu_{0}(\mathcal{X})=1$. In much of the literature, the measurement operator involves solving a PDE, in which case care must be taken when choosing the prior. The advantage of our approach is that the physics is encoded into the prior, rather than the likelihood. For the Brownian bridge, $\mu_{0}(L^{2}(\Omega))=1$, so the only requirement is that $R$ acts on $L^{2}(\Omega)$, a fairly trivial assumption.
To identify the MAP estimate of the posterior, we follow the work laid out in [dashti2013map]. The MAP estimate is identified through the Onsager-Machlup functional [durr1978onsager, fujita1982onsager]. This is the functional $I:H_{k}\to\mathbb{R}$ such that
$\lim_{\varepsilon\to 0}\frac{\mu^{y}(B(u_{2};\varepsilon))}{\mu^{y}(B(u_{1};\varepsilon))}=\exp\{I(u_{1})-I(u_{2})\},$
where $B(u_{i};\varepsilon)$ is the open ball in $L^{2}(\Omega)$ centered at $u_{i}$ with radius $\varepsilon$. For fixed $u_{1}$, any function $u_{2}$ which minimizes the Onsager-Machlup functional can be taken as the MAP estimate. For our specific problem, the Onsager-Machlup functional is
$I(u)=\left\{\begin{split}&\Phi(u)+\frac{1}{2}\|u-u^{0}\|^{2}_{H_{k}},\qquad\mathrm{if}\>u-u^{0}\in H_{k}\\ &+\infty,\qquad\mathrm{otherwise},\end{split}\right$. (3.6)
as shown in [dashti2013map, Theorem 3.2]. So, any MAP estimate of $\mu^{y}$ will live in the Cameron-Martin space (which is also $H_{k}$) of $\mu_{0}$. Further, if $\Phi$ is linear in $u$ this MAP estimate is unique.
Equation 3.6 is a natural candidate to build the variational problem for a general measurement operator. Adjusting the notation a bit to match the form of the linear case, the MAP estimate of the Gaussian measure solves the problem
$\hat{u}=\operatorname*{arg\,min}_{u\in H_{k}}\frac{1}{2}\|\Gamma^{-1}(y-R(u))\|^{2}+r\int_{\Omega}\frac{1}{2}\|\nabla u\|^{2}+qu\>d\Omega,$ (3.7)
which follows from Proposition 3.2. If the data are collected according to eq. (2.3), then it is easy to verify that eq. (3.7) reduces to the original variational problem.